Assume that the fixed cost of selling marketable securities is $10 per transaction and the interest rate on marketable securities is 8% per year. The company estimates that it will make cash payments of 12,500,000 per quarter.

Requires:
1. Optimal transaction size
2. average cash balance
3. the number of times(during the year) the company has to convert marketable securities to cash
4. total cost of converting marketable securities to cash
5. total carrying cost of cash
Answer #1
1] Optimal transaction size [Baumol Model] is given by:
C = (2A*F/O)^0.5, where
C = Optimum transaction size
A = Annual requirement of cash
F = Fixed converstion cost per transaction'
C = Opportunity cost of holding cash
Substituting available values in the above formula:
C = (2*12500000*4*10/0.08)^0.5 = $   1,11,803
2] Average cash balance = 111803/2 = $       55,902
3] Number of times conversion = 12500000*4/111803 = 447 times
4] Total cost of conversion = 447*10 = $         4,472
5) Total carrying cost = (111803/2)*8% = $         4,472